OpenStudy (anonymous):

Would someone please explain to me how to solve: (x+6)^2=20 1/2(x-5)-1=2/3(x-6) y(3y+7)=0 Thank you so much for the help

OpenStudy (anonymous):

\[\sqrt{(x+6)^{2}} = \sqrt{20}\] \[(x + 6) = \sqrt{4}\sqrt{5}\] \[x = -6 + 2\sqrt{5}\]

OpenStudy (anonymous):

Thank you so much, it looks clearer now :)

OpenStudy (anonymous):

Don't forget that the square roots are really \[\pm \sqrt{x}\] not just the positive root.

OpenStudy (anonymous):

alright, so would that answer be written as \[\pm x=2\sqrt{5}\] ?

OpenStudy (anonymous):

*\[\pm x = -6 + 2\sqrt{5}\]

OpenStudy (anonymous):

It would normally be written with x isolated like\[x = -6 \pm 2\sqrt{5}\]